On January 1, 2018, the Highlands Company began construction on a new manufacturing facility for its own use. The building was completed in 2019. The company borrowed $1,700,000 at 7% on January 1 to help finance the construction, in addition to the construction loan, Highlands had the following debt outstanding throughout 2018: $7,000,000, 12t bonds $3,000,000, 78 long-term note Construction expenditures incurred during 2018 were as follows: January 1 March 31 June 30 September 30 December 31 $ 740,000 1,340,000 968,000 740,000 540,000 Required: Calculate the amount of interest capitalized for 2018 using the specific interest method. (Do not round the intermediate calculations Round your percentage answers to i decimal place (i.e. 0.123 should be entered as 12.3%).) Date Wei January 1 March 31 June 30 September 30 December 31 Accumulated expenditure
Answer #1
Date expenditure Weight Average
January 1 $740,000 × 12/12 = $740,000
March 31 $1,340,000 × 9/12 = $10,05,000
June 30 $968,000 × 6/12 = $4,84,000
September 30 $740,000 × 3/12 = $1,85,000
December 31 $540,000 × 0/12 = 0
Accumulated expenditure $43,28,000 $24,14,000
Average Interest rate Capitalized interest
Average accumulated expenditure $24,14,000
Construction loan $17,00,000 × 7% = $119,000
Other loan(not construction) $7,14,000 × 10.5% = $74,970
$1,93,970

Explanation:-

Weighted-average rate of all other debt:-

$7,000,000 × 12% = $840,000

$3,000,000 × 7% = $210,000

$100,00,000 $10,50,000

= 10,50,000/$100,00,000 = 10.5%